RC meeting and Run Plan, Wednesday Sept. 27, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time: Monday - Friday at 3:30pm

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


Hall A call-in instructions:

  • Call RC first.
  • Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
  • After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

Hall Access procedure:

  • Access check list (binder on the counting room desk) needs to be filled out before making an access.
  • Scan (take pictures with phone) and post it to the logbook.
  • Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

New instruction for shift crew when changing target:

  • Beam restoration instructions have been updated. Printed version is in the counting room.
  • When we change target, make sure the SOFs of Target BPMs are set to the correct values.
  • SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)

Status Update

Overall Summary

Gordon Cates informed that target group preformed the magnetic field measurement inside the target enclosure during the maintenance day yesterday, and found no significant changes than before.

After the target group work, we closed the hall for beam permit around 6 pm. As PD mentioned before, MCC took couple of hours to recover beam after the maintenance. We got harp scans around the late swing shift yesterday, and beam recovery procedure was followed to send beam to the He3 target. We continued taking production data on polarized He3 (~40% polarization) @35uA beam current. No major issue during the OWL shift [except GEMs HV tripped and gap appeared in the occupancy plot].

In the PD meeting this morning, Jay mentioned that MCC found a solution to operate in higher power. So, he suggested MCC can try to send higher current to both halls A and C, if halls ready. During the RC meeting on Friday, we decided to go 40uA if we have polarization >35%. So I informed spokesperson regarding the increase of beam current upto 40uA. However around 09:05 am, PSS B-chain dropped BYI again and we don't have beam after that. We will ask MCC to send 40 uA when we get beam back to the hall.

Accelerator status

  • Issue with PSS in BSY again
  • No maintenance day in white board (until 10/10/2023)

Data collected :

  • Accumulated corrected charge collected: 0.434 C [during owl shift today] + 0 [yesterday day+swing shift]

Hall A Subsystems Status


  • Target polarization: 42% [12:20 pm]

Hall access request

Shift schedule

  • Still need to fill open slots
  • Open slots on Saturday 30th: DAY (TO) and SWING (SL).


Please sign up for shifts if there are any open!

Short Term Run plan

  • Beam recovery to the hall
  • Production run on polarized He3 target @40uA beam current (increase from 35uA to 40uA)

==> List of items in case He3 production running isn’t ready:

  1. Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
  2. BigBite Sieve slit => only during day (Techs needed)
    1. with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
    2. 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
  3. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)